OpenStudy (anonymous):

If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4. ?

OpenStudy (anonymous):

@esamalaa

OpenStudy (jdoe0001):

h = -16t2 + 64t + 4. <--- notice is a parabola, leading coefficient is negative, thus  so set y = 0 and solve for "x" \(\bf 0 = -16t^2 + 64t + 4\)

OpenStudy (jdoe0001):

"t" rather =)

OpenStudy (anonymous):

that's the prob.. I need help.

OpenStudy (jdoe0001):

\(\bf h= -16t^2 + 64t + 4\implies 0= -16t^2 + 64t + 4\\ \quad \\ \implies 0= -4(4t^2 - 16t - 1)\implies 0=(4t^2 - 16t - 1)\) so there, just factor it out

OpenStudy (anonymous):

is it 2(plus or minus) sqrt17/2?

OpenStudy (jdoe0001):

hmm dunno what factors did you get?

OpenStudy (jdoe0001):

I'd think you'd need to use the quadratic formula for this one, it won't factor out to integers

OpenStudy (jdoe0001):

\(\bf \qquad \qquad \qquad 0=({\color{blue}{ 4}}t^2 {\color{red}{ -16}} t {\color{green}{ -1}})\\ \textit{quadratic formula}\\ y={\color{blue}{ a}}x^2+{\color{red}{ b}}x+{\color{green}{ c}} \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)